OpenStudy (anonymous):

Can someone show me how to do: (8x + 2y)(5x - 4y)?

OpenStudy (anonymous):

I WILL!

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

okay so first what we wanna do is 8x*5x

OpenStudy (anonymous):

40x.

OpenStudy (anonymous):

Then 8x times -4y, correct?

OpenStudy (anonymous):

yes but instead of 40x it would be \[40x^2\]

OpenStudy (anonymous):

Oh! Okay. That's probably why I got it wrong. 8x times -4y equals -32xy, then you do 2y times 5x, which equals 10xy. Lastly, 2y times - 4y which equals - 8y. Then you combine terms.

OpenStudy (anonymous):

yes but remember the 8y should be\[8y^2\] because when you multiply two variables that are the same, you give it an exponent... so if it was x*x*x it would be \[x^3\]

OpenStudy (anonymous):

so you end with \[40x^2-22xy+8y^2\]

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

Could you help me with one more problem? If you don't mind.

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

Thank you. The problem is: (3x^2 + 2) + 4(x - 2)^2.

OpenStudy (anonymous):

I know first you do: (x - 2)^2 = (x - 2)(x - 2) = x^2 - 2x - 2x + 4.

OpenStudy (anonymous):

Simplified is x^2 - 4x^2 + 4.

OpenStudy (anonymous):

 this is what you should end with

OpenStudy (anonymous):

Oh, okay. Thank you!

OpenStudy (anonymous):

or at least im possative.... your explanation for simplifying is making me question myself

OpenStudy (anonymous):

nevermind im right just had to make sure

OpenStudy (anonymous):

Thanks again.

OpenStudy (anonymous):

okay wait I didn't fallow order of operations... the answer should be 7x^2-16x+18

OpenStudy (anonymous):

you were correct with the simplified x^2-4x+4 then add that o 3x^2 + 2 you get \[7x^2-16x+18\] sorry about that truthfully sorry

OpenStudy (anonymous):

No problem, thank you for your help.